OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

7 sin^2(x)-14sin(x)-7=0 , devide by seven sin^2(x)-2sin(x)+1=0 (sin(x)-1)^2=0 sin(x)=1 then x=pi/2 + 2n pi

OpenStudy (anonymous):

That's all you have to do?

OpenStudy (anonymous):

lol yes that's all, I am afraid you liked doing that :P